# SpeechJudge: Towards Human-Level Judgment for Speech Naturalness

### v1

@@ 1 Introduction
The collection and integration of human feedback corpora for model alignment has become a critical stage in the development of modern large-scale generative models, proving indispensable in domains such as text [44, 37, 3], image [54, 30], and video generation [55, 33].
@@
In the field of speech synthesis, naturalness has long been a cornerstone subjective metric for quality assessment [24, 2, 13, 56, 28], representing one of the most general-purpose indicators of performance [47, 45]. Prior research has explored automated speech assessment through MOS predictors [41, 21] and constructed the human feedback corpora for specific attributes like the low-level acoustic quality [50]. However, a large-scale human feedback corpus centered on the holistic quality of naturalness—and a corresponding reward model trained to capture these preferences—remains a notably underexplored area. To fill this void, this paper focuses on the dimension of speech naturalness and present a three-part contribution:
@@
1. A Large-scale Human Feedback Dataset: SpeechJudge-Data. We recruit human annotators to provide feedback on synthesized speeches, with a focus on assessing two fundamental speech aspects: intelligibility and naturalness. For data synthesis, we employ a diverse set of advanced, open-source zero-shot TTS models with varying architectures (such as CosyVoice2 [14], Ints [61], F5-TTS [7], and MaskGCT [52]) to produce the compared speech pairs. We prepare speech references in both regular and expressive styles, construct multilingual target texts, and cover both monolingual and cross-lingual synthesis scenarios to ensure data diversity (Section 3.1). We instruct human annotators to perform two tasks based on a speech pair (Figure 1): (a) pointwise annotation of text accuracy to assess intelligibility, and (b) pairwise preference annotation to judge relative speech naturalness. This extensive effort, involving 69 labelers over two months, results in 99K annotated pairs, with each pair receiving an average of 2.49 annotations from different labelers. We believe the SpeechJudge-Data can serve as a valuable corpus for alignment research in speech synthesis (e.g., DPO alignment [39] or reward modeling [44, 37, 3] in Section 5).
@@
2. An Evaluation Benchmark for Speech Naturalness Judgment: SpeechJudge-Eval. We design a dedicated evaluation benchmark for the task of speech naturalness judgment. The task is structured as follows: given a target text and two corresponding speech samples, a model needs to judge which one is more natural. To construct the evaluation set, we select a subset from the SpeechJudge-Data where human annotators demonstrated high inter-annotator agreement, ensuring a high-quality ground truth. We assess the naturalness judgment capabilities of a wide range of metrics and models, including Word Error Rate (WER) [38, 16], Fréchet Audio Distance (FAD) [27], MOS predictors [41, 40, 48], Deepfake Detectors [25, 49], and AudioLLMs [56, 28, 53, 8, 22]. Our evaluations reveal that even the most capable model—specifically, Gemini-2.5-Flash [8] in our experiments—achieved less than 70% agreement with human preferences. This finding highlights a significant performance gap and underscores the substantial room for research and improvement in automated speech naturalness judgment.
@@
3. A Generative Reward Model for Speech Naturalness: SpeechJudge-GRM. To develop a reward model that more effectively captures human preferences, we develop SpeechJudge-GRM, a generative reward model (GRM) [59, 34] trained on the SpeechJudge-Data. Specifically, we base our model on Qwen2.5-Omni-7B [56] and design a two-stage post-training process. During the first stage, we perform SFT as the “cold start” to improve the model’s instruction-following and rationale-based reasoning capabilities. To achieve this, we leverage Gemini-2.5-Flash [8] to generate Chain-of-Thought (CoT) data for speech naturalness judgment task. In the second stage, we focus on more challenging cases of SpeechJudge-Data, which we define as instances where Gemini-2.5-Flash fails to make the correct judgment. Treating the human-annotated labels as the verifiable reward [11, 34], we apply the GRPO-based RL stage [42]. Our experiments demonstrate that when trained on the same data, SpeechJudge-GRM significantly outperformed the classic Bradley-Terry reward model (BTRM) [4, 39], achieving a higher accuracy in predicting human preferences (77.2% for SpeechJudge-GRM vs. 72.7% for SpeechJudge-BTRM, Table 3). Besides, SpeechJudge-GRM also supports inference-time scaling and offers explainability through its CoT outputs. Furthermore, SpeechJudge-GRM can also be employed as an objective naturalness metric for sample selection (Figure 5) or as a reward function in RL algorithms to enhance the quality of existing speech generation models (Figure 6).
@@ 2 Related Work
Human Alignment for Speech Generation Aligning generative models with human feedback has proven crucial, a process also known as RLHF in LLMs [37, 3]. In the vision domain, many similar human preference datasets exist, such as Pick-a-Pic [30], ImageReward [54], and VideoReward [33]. The speech synthesis field, pioneering efforts to construct human corpora involved MOS datasets [41, 21]. However, these datasets often did not use advanced TTS models for data generation, provided only the pointwise labels rather than the direct pairwise human preference, and were limited in scale. More recently, efforts have focused on building human feedback corpora centered on specific speech attributes, such as low-level acoustic quality [50], intelligibility [61], or the instruction-following capabilities of spoken dialogue systems [23, 17]. Despite this progress, a large-scale human feedback corpus built specifically around naturalness—one of the most general-purpose and fundamental metrics for speech synthesis [47, 45]—has remained a critical missing piece.
AudioLLM as a Judge Using LLMs as automated quality evaluators is a prominent topic in the textual LLM field, popularized by the “LLM-as-a-judge” paradigm [64]. This approach has recently been extended to the speech domain. A concurrent work, AudioJudge [36], evaluates the capabilities and limitations of using AudioLLMs for speech quality assessment and paralinguistic understanding by prompt engineering. Furthermore, many studies have focused on fine-tuning AudioLLMs to elicit their understanding capabilities for specific tasks. Examples include discriminating the human-likeness of audio [51], understanding low-level acoustic qualities [5, 50], and enhancing the assessment of instruction-following in spoken dialogue systems [23, 17]. However, how to improve the ability of AudioLLMs to understand and judge speech naturalness, and how to use the quality assessment capabilities of AudioLLMs as a reward to improve the post-training of speech generation models themselves, remain significantly underexplored.
@@ 3.1 Dataset Construction
Model Selection For $\mathcal{M}_{tts}$, we select the following six models of three architectures to enrich the distribution of the synthetic data (Figure 2(a)): (1) AR-based: ARS [52], CosyVoice2 [14], CosyVoice2-INTP [61], and Ints-INTP [61]. The latter two are released by Zhang et al. [61] as intelligibility-enhanced models. (2) FM-based: F5-TTS. (3) MGM-based: MaskGCT [52].
Prompt Construction To build diverse prompts $(a_{ref},t)$ for TTS, for $a_{ref}$, we adopt both regular and expressive speech samples. The regular samples are randomly selected from the Emilia-Large dataset [19]. The expressive samples are sourced from corpora rich in paralinguistics, including the emotional corpora: ParaSpeechCaps [9], the accented corpora: L2-Arctic [63] and KeSpeech [46], the whisper samples from an in-house corpus, and the character voices from video games Genshin Impact [43]. We display the detailed distribution of speech references in Figure 2(b).
The target text $t$ paired with each $a_{ref}$ is constructed as follows: For regular $a_{ref}$ samples, we randomly sample transcriptions from the Emilia-Large dataset [19]. These are then refined using DeepSeek-V3 [10] to correct typos and normalize punctuations. For expressive $a_{ref}$ samples, we instruct DeepSeek-V3 to generate several scripts in different writing styles, tailored to the topic of $a_{ref}$ (see Appendix 7.1 for more details). The languages of the target texts included Chinese (zh), English (en), and Chinese-English code-switching (mixed). For the combinations $(a_{ref},t)$, we include both monolingual settings (en2en and zh2zh) and cross-lingual settings (zh2en, en2zh, zh2mixed, and en2mixed), where zh2en denotes Chinese $a_{ref}$ with English $t$, and similarly for others. The distribution of the language settings of $(a_{ref},t)$ is shown in Figure 2(c).
@@ 3.2 Human Annotation
Given a sample $(t,a_{1},a_{2})$, human annotators are instructed to perform both pointwise intelligibility and pairwise naturalness annotations (Figure 1). For intelligibility, annotators perform a binary classification to determine whether the speech ($a_{1}$ and $a_{2}$) accurately reads the text $t$ without any content insertion, omission, or mispronunciation. For naturalness, they perform a five-scale Comparative Mean Opinion Score (CMOS) annotation to determine which of the two audio clips ($a_{1}$ or $a_{2}$) sounds more natural and human-like. We recruit human annotators and provide them with training. The detailed annotation guidelines are provided in Appendix 8.
@@ 4 SpeechJudge-Eval
To evaluate speech naturalness, existing studies typically organize their own listening tests, which often have inconsistent settings across different papers [24, 2, 13, 56, 28]. Alternatively, previous researchers use proxy MOS predictors, such as UTMOS [41], as an objective metric. However, it remains an underexplored problem whether these metrics can accurately judge the naturalness of more advanced speech generation models [52, 14, 7, 61] and align with human preferences. Motivated by this, we construct a benchmark, SpeechJudge-Eval, specifically for the speech naturalness judgment task.
@@ 4.2 Benchmark for Different Models
    Objective metrics, such as WER [38, 16], SIM [6], and FAD [27] in audio generation tasks. We assume that a better value of these metrics (e.g., lower for WER and FAD; higher for SIM) indicates better naturalness.
@@
    MOS Predictors, including DNSMOS [40], UTMOS [41], and predictors from audiobox-aesthetics (CE, CU, PC, and PQ) [48]. We assume that a higher MOS score corresponds to better naturalness.
@@
    Deepfake detectors, which are typically pre-trained on a binary classification task to predict whether an audio is fake or not [25, 49]. We assume that an audio with a lower fake probability should have better naturalness.
@@
    AudioLLMs, which are employed to test their speech naturalness understanding capabilities in a zero-shot manner [^2] We assume that the adopted AudioLLMs have not been directly trained on the speech naturalness judgment task. Their performance on this benchmark is therefore considered a zero-shot capability.. We include the open-source Phi-4-Multimodal [1], Qwen2.5-Omni [56], Kimi-Audio [28], Gemma-3n [26], Voxtral [32], MiDashengLM [12], Mimo-Audio [53], and the closed-source Gemini-2.5 [8] and GPT-4o [22]. We use the plain prompt of Table 1 to instruct the model to pairwise score the naturalness of two audios. We use their grading to determine the naturalness preference.
The performance of different models on SpeechJudge-Eval is presented in Table 2. A key observation is that speech naturalness judgment is a highly challenging task. The leading model, Gemini-2.5-Flash, still only achieves less than 70% agreement with human preferences. When comparing different models, we find that: (1) common objective metrics and MOS predictors show only a weak correlation with human preferences, often achieving less than 60% accuracy and sometimes performing at the level of a random guess (around 50%). (2) While deepfake detectors are highly effective at distinguishing between machine-generated and human-recorded speech [49, 25], their ability to do so is not well-aligned with the naturalness objective when comparing two generated samples. (3) AudioLLMs demonstrate significant potential for this task. While some models, such as Gemma-3n and GPT-4o mini Audio, perform at a chance level, a number of others achieve an accuracy exceeding 60%. This promising performance motivates us to further leverage these AudioLLMs for the design of a reward model for speech naturalness.
@@ 5 SpeechJudge-GRM
Based on the proposed SpeechJudge-Data, we further explore how to train a reward model capable of accurately capturing human preferences. Specifically, we propose SpeechJudge-GRM, where we leverage the inherent audio understanding capabilities of AudioLLMs (specifically, Qwen2.5-Omni-7B [56]) to elicit their speech naturalness judgment capability. Compared to the classic BTRM [4], the key strengths of GRM are its ability to enable Chain-of-Thought (CoT) reasoning and its support for test-time computation via majority voting, which ultimately leads to improved preference judgment performance [59].

### v2

@@ 1 Introduction
The collection and integration of human feedback corpora for model alignment has become a critical stage in the development of modern large-scale generative models, proving indispensable in domains such as text [openai-rm, instructgpt, RLHF-anthoropic], image [imagereward, pick-a-pic], and video generation [visionreward, videoreward].
@@
In the field of speech synthesis, naturalness has long been a cornerstone subjective metric for quality assessment [naturalspeech3, seedtts, cosyvoice, qwen2.5-omni, kimi-audio], representing one of the most general-purpose indicators of performance [tts-book-paul-taylor, tts-book-tanxu]. Prior research has explored automated speech assessment through MOS predictors [utmos, voicemos-challenge-2024] and constructed the human feedback corpora for specific attributes like the low-level acoustic quality [qualispeech]. However, a large-scale human feedback corpus centered on the holistic quality of naturalness—and a corresponding reward model trained to capture these preferences—remains a notably underexplored area. To fill this void, this paper focuses on the dimension of speech naturalness and present a three-part contribution:
@@
1. A Large-scale Human Feedback Dataset: SpeechJudge-Data. We recruit human annotators to provide feedback on synthesized speeches, with a focus on assessing two fundamental speech aspects: intelligibility and naturalness. For data synthesis, we employ a diverse set of advanced, open-source zero-shot TTS models with varying architectures (such as CosyVoice2 [cosyvoice2], Ints [intp], F5-TTS [f5tts], and MaskGCT [maskgct]) to produce the compared speech pairs. We prepare speech references in both regular and expressive styles, construct multilingual target texts, and cover both monolingual and cross-lingual synthesis scenarios to ensure data diversity (Section 3.1). We instruct human annotators to perform two tasks based on a speech pair (Figure 1): (a) pointwise annotation of text accuracy to assess intelligibility, and (b) pairwise preference annotation to judge relative speech naturalness. This extensive effort, involving 69 labelers over two months, results in 99K annotated pairs, with each pair receiving an average of 2.49 annotations from different labelers. We believe the SpeechJudge-Data can serve as a valuable corpus for alignment research in speech synthesis (e.g., DPO alignment [dpo] or reward modeling [openai-rm, instructgpt, RLHF-anthoropic] in Section 5).
@@
2. An Evaluation Benchmark for Speech Naturalness Judgment: SpeechJudge-Eval. We design a dedicated evaluation benchmark for the task of speech naturalness judgment. The task is structured as follows: given a target text and two corresponding speech samples, a model needs to judge which one is more natural. To construct the evaluation set, we select a subset from the SpeechJudge-Data where human annotators demonstrated high inter-annotator agreement, ensuring a high-quality ground truth. We assess the naturalness judgment capabilities of a wide range of metrics and models, including Word Error Rate (WER) [whisper, funasr], Fréchet Audio Distance (FAD) [fad], MOS predictors [utmos, dnsmos, meta-audiobox-aesthetics], Deepfake Detectors [aasist, audio-deepfake-verification], and AudioLLMs [qwen2.5-omni, kimi-audio, mimoaudio, gemini2.5, gpt4o]. Our evaluations reveal that even the most capable model—specifically, Gemini-2.5-Flash [gemini2.5] in our experiments—achieved less than 70% agreement with human preferences. This finding highlights a significant performance gap and underscores the substantial room for research and improvement in automated speech naturalness judgment.
@@
3. A Generative Reward Model for Speech Naturalness: SpeechJudge-GRM. To develop a reward model that more effectively captures human preferences, we develop SpeechJudge-GRM, a generative reward model (GRM) [google-grm, deepseek-grm] trained on the SpeechJudge-Data. Specifically, we base our model on Qwen2.5-Omni-7B [qwen2.5-omni] and design a two-stage post-training process. During the first stage, we perform Supervised Fine-Tuning (SFT) as the “cold start” to improve the model’s instruction-following and rationale-based reasoning capabilities. To achieve this, we leverage Gemini-2.5-Flash [gemini2.5] to generate Chain-of-Thought (CoT) data for speech naturalness judgment task. In the second stage, we focus on more challenging cases of SpeechJudge-Data, which we define as instances where Gemini-2.5-Flash fails to make the correct judgment. Treating the human-annotated labels as the verifiable reward [deepseek-r1, deepseek-grm], we apply the GRPO-based Reinforcement Learning (RL) stage [grpo]. Our experiments demonstrate that when trained on the same data, SpeechJudge-GRM significantly outperformed the classic Bradley-Terry reward model (BTRM) [bt-rm, dpo], achieving a higher accuracy in predicting human preferences (77.2% for SpeechJudge-GRM vs. 72.7% for SpeechJudge-BTRM, Table 3). Besides, SpeechJudge-GRM also supports inference-time scaling and offers explainability through its CoT outputs. Furthermore, SpeechJudge-GRM can also be employed as an objective naturalness metric for sample selection (Figure 5) or as a reward function in RL algorithms to enhance the quality of existing speech generation models (Figure 6).
We will release all resources at https://github.com/AmphionTeam/SpeechJudge to facilitate future research in human-aligned speech synthesis. Audio samples are available at https://speechjudge.github.io/.
@@ 2 Related Work
Human Alignment for Speech Generation Aligning generative models with human feedback has proven crucial, a process also known as RLHF in LLMs [instructgpt, RLHF-anthoropic]. In the vision domain, many similar human preference datasets exist, such as Pick-a-Pic [pick-a-pic], ImageReward [imagereward], and VideoReward [videoreward]. The speech synthesis field, pioneering efforts to construct human corpora involved MOS datasets [utmos, voicemos-challenge-2024]. However, these datasets often did not use advanced TTS models for data generation, provided only the pointwise labels rather than the direct pairwise human preference, and were limited in scale. More recently, efforts have focused on building human feedback corpora centered on specific speech attributes, such as low-level acoustic quality [qualispeech], intelligibility [intp], or the instruction-following capabilities of spoken dialogue systems [wavreward, sagelm]. Despite this progress, a large-scale human feedback corpus built specifically around naturalness—one of the most general-purpose and fundamental metrics for speech synthesis [tts-book-paul-taylor, tts-book-tanxu]—has remained a critical missing piece.
AudioLLM as a Judge Using LLMs as automated quality evaluators is a prominent topic in the textual LLM field, popularized by the “LLM-as-a-judge” paradigm [llm-as-a-judge]. This idea has recently been extended to the audio domain. A concurrent work, AudioJudge [audiojudge], evaluates the capabilities and limitations of using AudioLLMs for speech quality assessment and paralinguistic understanding via prompt engineering. Furthermore, many studies have focused on fine-tuning AudioLLMs to better expose their understanding capabilities for specific tasks, such as discriminating the human-likeness of audio [audio-turing-test], modeling low-level acoustic qualities [chenchen-quality-audiollm, qualispeech], unifying multiple speech quality evaluation tasks into a single AudioLLM [sq-llm], and enhancing the assessment of instruction-following in spoken dialogue systems [wavreward, sagelm]. However, how to improve the ability of AudioLLMs to understand and judge speech naturalness, and how to use their quality-assessment capabilities as a reward to improve the post-training of speech generation models themselves, remain significantly underexplored.
@@ 3.1 Dataset Construction
Model Selection For $\mathcal{M}_{tts}$, we select the following six models of three architectures to enrich the distribution of the synthetic data (Figure 2(a)): (1) AR-based: ARS [maskgct], CosyVoice2 [cosyvoice2], CosyVoice2-INTP [intp], and Ints-INTP [intp]. The latter two are released by intp as intelligibility-enhanced models. (2) FM-based: F5-TTS. (3) MGM-based: MaskGCT [maskgct].
Prompt Construction To build diverse prompts $(a_{ref},t)$ for TTS, for $a_{ref}$, we adopt both regular and expressive speech samples. The regular samples are randomly selected from the Emilia-Large dataset [emilia-large]. The expressive samples are sourced from corpora rich in paralinguistics, including the emotional corpora: ParaSpeechCaps [paraspeechcaps], the accented corpora: L2-Arctic [l2arctic] and KeSpeech [kespeech], the whisper samples from an in-house corpus, and the character voices from video games Genshin Impact [genshin]. We display the detailed distribution of speech references in Figure 2(b).
The target text $t$ paired with each $a_{ref}$ is constructed as follows: For regular $a_{ref}$ samples, we randomly sample transcriptions from the Emilia-Large dataset [emilia-large]. These are then refined using DeepSeek-V3 [deepseek-v3] to correct typos and normalize punctuations. For expressive $a_{ref}$ samples, we instruct DeepSeek-V3 to generate several scripts in different writing styles, tailored to the topic of $a_{ref}$ (see Appendix 8.1 for more details). The languages of the target texts included Chinese (zh), English (en), and Chinese-English code-switching (mixed). For the combinations $(a_{ref},t)$, we include both monolingual settings (en2en and zh2zh) and cross-lingual settings (zh2en, en2zh, zh2mixed, and en2mixed), where zh2en denotes Chinese $a_{ref}$ with English $t$, and similarly for others. The distribution of the language settings of $(a_{ref},t)$ is shown in Figure 2(c).
@@ 3.2 Human Annotation
Given a sample $(t,a_{1},a_{2})$, human annotators are instructed to perform both pointwise intelligibility and pairwise naturalness annotations (Figure 1). For intelligibility, annotators perform a binary classification to determine whether the speech ($a_{1}$ and $a_{2}$) accurately reads the text $t$ without any content insertion, omission, or mispronunciation. For naturalness, they perform a five-scale Comparative Mean Opinion Score (CMOS) annotation to determine which of the two audio clips ($a_{1}$ or $a_{2}$) sounds more natural and human-like.
We recruited professional annotators from a specialized data annotation firm in China and provided them with training for speech naturalness judgement. All annotators assigned to Chinese data were native speakers. For the English and code-switching datasets, annotators were required to have a proficiency level equivalent to at least CET-6. All personnel underwent standardized training based on a detailed annotation manual. Initially, we conducted a pilot study among researchers to refine the guidelines for clarity and unambiguity. To ensure annotation quality, each sample $(t,a_{1},a_{2})$ was independently annotated by two individuals. A third annotator was introduced if any disagreements. The detailed annotation guidelines are provided in Appendix 9.
@@ 4 SpeechJudge-Eval
To evaluate speech naturalness, existing studies typically organize their own listening tests, which often have inconsistent settings across different papers [naturalspeech3, seedtts, cosyvoice, qwen2.5-omni, kimi-audio]. Alternatively, previous researchers use proxy MOS predictors, such as UTMOS [utmos], as an objective metric. However, it remains an underexplored problem whether these metrics can accurately judge the naturalness of more advanced speech generation models [maskgct, cosyvoice2, f5tts, intp] and align with human preferences. Motivated by this, we construct a benchmark, SpeechJudge-Eval, specifically for the speech naturalness judgment task.
@@ 4.2 Benchmark for Different Models
    Objective metrics, such as WER [whisper, funasr], SIM [wavlm], and FAD [fad] in audio generation tasks. We assume that a better value of these metrics (e.g., lower for WER and FAD; higher for SIM) indicates better naturalness.
@@
    MOS Predictors, including DNSMOS [dnsmos], UTMOS [utmos], and predictors from audiobox-aesthetics (CE, CU, PC, and PQ) [meta-audiobox-aesthetics]. We assume that a higher MOS score corresponds to better naturalness.
@@
    Deepfake detectors, which are typically pre-trained on a binary classification task to predict whether an audio is fake or not [aasist, audio-deepfake-verification]. We assume that an audio with a lower fake probability should have better naturalness.
@@
    AudioLLMs, which are employed to test their speech naturalness understanding capabilities in a zero-shot manner [^2] We assume that the adopted AudioLLMs have not been directly trained on the speech naturalness judgment task. Their performance on this benchmark is therefore considered a zero-shot capability.. We include the open-source Phi-4-Multimodal [phi4-mm], Qwen2.5-Omni [qwen2.5-omni], Kimi-Audio [kimi-audio], Gemma-3n [gemma3], Voxtral [voxtral], MiDashengLM [midashenglm], Mimo-Audio [mimoaudio], and the closed-source Gemini-2.5 [gemini2.5] and GPT-4o [gpt4o]. We use the plain prompt of Table 1 to instruct the model to pairwise score the naturalness of two audios. We use their grading to determine the naturalness preference.
The performance of different models on SpeechJudge-Eval is presented in Table 2. A key observation is that speech naturalness judgment is a highly challenging task. The leading model, Gemini-2.5-Flash, still only achieves less than 70% agreement with human preferences. When comparing different models, we find that: (1) common objective metrics and MOS predictors show only a weak correlation with human preferences, often achieving less than 60% accuracy and sometimes performing at the level of a random guess (around 50%). (2) While deepfake detectors are highly effective at distinguishing between machine-generated and human-recorded speech [audio-deepfake-verification, aasist], their ability to do so is not well-aligned with the naturalness objective when comparing two generated samples. (3) AudioLLMs demonstrate significant potential for this task. While some models, such as Gemma-3n and GPT-4o mini Audio, perform at a chance level, a number of others achieve an accuracy exceeding 60%. This promising performance motivates us to further leverage these AudioLLMs for the design of a reward model for speech naturalness.
@@ 5 SpeechJudge-GRM
Based on the proposed SpeechJudge-Data, we further explore how to train a reward model capable of accurately capturing human preferences. Specifically, we propose SpeechJudge-GRM, where we leverage the inherent audio understanding capabilities of AudioLLMs (specifically, Qwen2.5-Omni-7B [qwen2.5-omni]) to elicit their speech naturalness judgment capability. Compared to the classic BTRM [bt-rm], the key strengths of GRM are its ability to enable Chain-of-Thought (CoT) reasoning and its support for test-time computation via majority voting, which ultimately leads to improved preference judgment performance [google-grm].